OpenStudy (anonymous):

Factor completely if possible. -2x^2 -x+36

OpenStudy (jhannybean):

\[-2x^2 -x+36\]\[x^2 -x -72\]\[(x-9)(x+8)\]\[\left(x-\frac{9}{-2}\right)\left(x+\frac{8}{-2}\right)\]if the fraction doesn't simplify, multiply the \(-2\) to the variable\[(-2x-9)(x-4)\]

OpenStudy (jhannybean):

What I did was I multiplied the \(-2\) to \(36\) to simplify my expression. then I found two numbers that multiply to give me -72 and add to give me -1. Because I multiplied -2 to simplify my equation, I had to divide by it to reinclude it into my factored form. Then I just simplified.